# VibraForge: A Scalable Prototyping Toolkit For Creating Spatialized Vibrotactile Feedback Systems

### v1

@@ 3.2.1. Vibrotactile Actuators
An X-axis, square-shaped, 9.5 mm x 9.5 mm LRA manufactured by the Jinlong Machinery and Electronics Company (Model L959535) (Machinery and Co., 2022) was chosen for the vibration unit. Its resonant frequency was 170 Hz. When powered by 0.9 Vrms, it had a maximum current of 145 mArms and an acceleration of 0.65±0.12 Grms. The other two actuators were a Z-axis LRA manufactured by Jinlong (Model G0832) (Machinery and Co., 2022) and a VCA manufactured by PUI Audio, Inc. (Model HD-VA3222) (PUI Audio, 2023). The Z-axis LRA was cylindrical-shaped, with an 8 mm diameter and a height of 3.2 mm. Its resonant frequency was 235 Hz. When powered by 1.8 Vrms, it had a maximum current of 90 mArms and an acceleration of 0.50±0.08 Grms. The VCA was 32 mm x 22 mm and had a frequency response range from 80 - 500 Hz. When driven by a 133 Hz 1.5 Vrms signal, it had a maximum acceleration of 2.52 Gp-p.
These actuators were chosen because they cover a variety of vibration directions and intensities, thus satisfying the tactile perception requirements of different body parts (i.e., spatial acuity (Elsayed et al., 2020) and perceived intensity (Verrillo et al., 1969)). Since they all share the same PCB schematics and similar enclosure structures, in the following sections we only describe the vibration unit designed for the X-axis LRA. The other designs can be seen in Figure 3.
@@ 3.2.2. Printed Circuit Board Design
Due to bandwidth limitations, as the number of actuators increases it becomes difficult to drive them using real-time PWM signals sent from a central unit. Therefore, we allocated an MCU to each vibration unit to drive each actuator. Each vibration unit had its own driver PCB to receive UART commands from a central board and provide fine-grained control over vibration characteristics. The driver PCB had a PIC16F18313 MCU, a DRV8837 H-bridge motor driver, two 4-pin JST-SH headers, two pogo pins, and supportive components (Figure 4). The MCU was used to transmit commands and generate waveforms. The JST headers were used as input and output for the power and signal connections. The number of connector pins was minimized to four (i.e., actuator power, MCU power, ground, and data). The pogo pins were mounted on the bottom to ensure direct contact with the actuator pads.
Figure 4. The circuit schematics and PCB board designs of vibration units.
@@
When the MCU received a ”start” signal, it sent a waveform to the H-bridge to modulate the actuator. We utilized the PWM generator inside the MCU for fine-grained control over the actuators’ intensities, frequencies, and waveforms. Because there is no address system for each unit, designers only need to load their program into the MCU once. They can then plug-and-play each unit as desired. The communication protocol is further described in Section 3.4.
@@
The 3D printed enclosures were designed to enclose each vibrotactile actuator and PCB board and provide a method to easily attach and detach each vibration unit to a garment (DR4). Similar to the size of a coin, the enclosure was compact and lightweight, enabling for a comfortable wearing experience (DR5).
The enclosure consisted of four parts: a cap, a PCB mount, an enclosure, and a bottom ring (Figure 3). The cap employed a screw-nut fastening system, which added top-down pressure to stabilize the PCB. The PCB mount offered protection for the PCB and aligned the pogo pins directly with the actuator pads. This improved the robustness of the vibration unit as no soldering was required during assembly. The enclosure served as a protective housing for all the parts. It had two holes for JST connectors, as well as two asymmetrically positioned vents for passive heat dissipation. The bottom ring was designed to be positioned beneath the garment to maintain direct contact with the skin without the need for any sewing (DR4).
The same design can be easily adapted to other vibration unit structures for other types of actuators. In Figure 3, other vibration units that share the same PCB schematics and slightly different enclosure designs are shown.
@@ 3.3. Control Unit
The control unit was designed to be fully portable (DR6). It consisted of a PCB extension board with an ESP32 microcontroller (MCU), two LiPo batteries, and a 3D-printed enclosure (Figure 5). Multiple versions of control units were designed to accommodate different projects, ranging from simple arm bands to more spatialized body displays. They varied in the (a) number of chains supported, (b) model of the ESP32 board, and (c) battery capacity.
@@
We designed a large control unit that supported up to 8 chains of vibration units and 16 units per chain, for a total of 128 vibration units.The extension board drew power from two serial 3.7 V 8200 mAh LiPo batteries. It stabilized the voltage at 5 V through two power regulators: a linear regulator that powered all the MCUs and a buck regulator that powered all actuators in the system. These separate power supplies ensured that MCUs operated on stable voltages without interference from the actuators’ power consumption. The extension board had eight JST connectors. Each connector was connected to a GPIO pin on the ESP32 MCU and drove a chain of vibration units. The MCU was an Adafruit Feather ESP32 V2222https://learn.adafruit.com/adafruit-esp32-feather-v2. It was responsible for converting external Bluetooth commands to UART commands and sending them to the corresponding vibration unit.
Since the number of required vibration units varies across different tasks, we also designed a small control unit with fewer supported chains. This compact unit design contained two 3.7 V 500 mAh LiPo batteries and an Adafruit QT Py ESP32-S3  [^3] https://learn.adafruit.com/adafruit-qt-py-esp32-s3/pinouts. This MCU had four general purpose GPIO pins that could be used for chain connections.
@@ 3.4. Chain-Connection Method and Data Communication
VibraForge differs from existing toolkits due to its use of a chain-connection method (Figure 6). This method enables the toolkit to address over 100 vibration units using a single control unit (DR1) and support the simultaneous control over the vibration characteristics of multiple actuators (DR2).
@@
As explained in Section 3.3, each GPIO pin on the ESP32 MCU was routed to a JST connector on the extension board. Because vibration units shared the same JST connectors, they could be connected in chains and receive commands through the same GPIO pin. There were two options for the data transmission design. The first option was to use data bus protocols such as CAN or I2C, where all the vibration units would receive the same command simultaneously and use the command’s address to determine whether the command was intended for itself. While this option would allow commands to be transmitted with a minimal latency, the vibration units would need to be programmed with different addresses prior to use, which would be burdensome when working with multiple units. Instead, we decided to send commands through each chain using a serial protocol, where each unit would receive a command, process it, and resend it to the next unit. While this may increase the overall latency, it allowed each unit to operate with a ”virtual address”, enabling independent plug-and-play control without additional programming.
Figure 7. The two-byte message bits definition and an example of how vibration commands are transmitted along a chain using the ”virtual address” system.
@@
Vibration commands were transmitted via a UART protocol at 115.2 kHz with parity checks (Figure 7). Each command consisted of two bytes: the first byte encoded the address and start/stop status of the target vibration unit, and the second byte detailed the vibration characteristics. Upon receiving the first byte, a vibration unit checked the address. A non-zero address indicated the command was for subsequent units, prompting the unit to decrement the address and forward the message. This process continued until the address reached zero and signified the targeted unit, which then acted based on the start/stop bit: ’1’ (START) awaited the second byte, while ’0’ (STOP) disabled its actuator. When a unit was ”signified”, it waited for the second-byte message, which contained details about vibration characteristics, including the 16 intensity levels (evenly distributed from 0% to 100% duty cycle) and 8 frequencies (i.e., 123, 145, 170, 200, 235, 275, 322, and 384 Hz). The frequency levels were chosen so that each level was 1.18 times higher than the previous level. This followed Weber’s factor for vibrotactile frequency just-noticeable differences (Pongrac, 2006).
@@ 3.5. Signal Segmentation Algorithm
To render complex waveforms such as audio signals on vibration units (DR2), we implemented a signal segmentation algorithm that converted high-frequency input signals to low-frequency serial commands (Figure 8).For instance, consider an input waveform of 44100 Hz audio signals that composed of 200 Hz and 5 Hz sine waves. High-frequency components (200 Hz) in the waveform were extracted using a short-time Fourier transform (STFT), which captured how the dominant frequency changed over time. Low-frequency components (5 Hz) were extracted using a Hilbert transform, which captured the envelope of the waveform. Both components were downsampled to 200 Hz and sent to vibration units. Since our frequency levels ranged from 123 Hz to 384 Hz, the segmentation threshold was set to 100 Hz. Frequency components above 100 Hz were rendered using frequency levels, while components below 100 Hz were rendered using intensity levels. The technical evaluations in Section 4 demonstrated that this algorithm could reliably render complex waveforms such as audio signals using serial command control.
@@ 3.6. GUI Editor for Designing Spatialized Patterns
We designed a GUI Editor (Figure 9) to enable designers to create spatialized vibration patterns (DR3). This editor differed from previous editors in that it combined the waveform editing from single-actuator design software with the canvas design used for multi-actuator design. The key components of the GUI Editor were:
@@
  • Waveform Library: The Waveform Library contained waveform templates such as oscillators and envelopes. Designers could use them to create a custom waveform and save it to the Waveform Library. In addition, designers could import waveforms in the JSON format from popular waveform editors such as Syntacts (Pezent et al., 2020), VIREO (Terenti and Vatavu, 2023), and Meta Haptics Studio (Meta Platforms, 2023a) so they could reuse previous designs (DR3).
  • Waveform Visualizer: The Waveform Visualizer served two purposes. When designers were creating a custom waveform, they could drag and drop waveform templates from the Waveform Library onto the Waveform Visualizer and it would show the combined waveform. When designers were composing vibration patterns for individual vibration units, the Waveform Visualizer also show the composed waveform for the selected unit.
  • Canvas: The Canvas enabled designers to drag-and-drop vibration units onto the canvas so they could be connected into chains that mimicked the physical layout of the vibration units. In addition, a ”Create New Chain” function enabled designers to generate vibration units in batches and arrange them using grid layouts. When clicking on a unit, its corresponding waveform would show up in the Waveform Visualizer.
  • Timeline: The Timeline served as a high-level summary of the vibration patterns. The blue blocks showed the range of time when the vibration unit was activated. By moving the slider left and right, the vibration units that were activated at that moment would be highlighted in the Canvas so designers could understand the vibration patterns over time. By clicking the ”Play” button, the waveform would be converted into vibration commands with timestamps and sent to the control unit via Bluetooth.
In a typical spatial pattern design workflow, designers would begin by creating a new design and specify the number of chains and actuators to be used in their system. The GUI would then generate a default Canvas with the designated number of actuators. Designers would then freely adjust the layout to match their physical device. Then, designers would assign waveforms to each vibration unit and specify the start and end time. Each waveform could be designed from scratch using the waveform templates in the Waveform Library, or imported from other software. To review the spatialized patterns, designers would click on any vibration unit on the Canvas and the Waveform Visualizer would show the corresponding waveform. Alternatively, designers could see a high-level summary of how the vibration patterns of multiple units would change over time by moving the Timeline slider.
Additionally, we also provided a Python server with VibraForge APIs, so that vibration units can be directly activated and controlled via code (DR3). Later in Section 5, Python servers are integrated into external programming environments (e.g., Unity) to trigger real-time feedback for applications.
@@ 4. Technical Evaluation
In this section, we summarize various performance metrics of the VibraForge toolkit to demonstrate how it fulfills the aforementioned design requirements. All tests below used vibration units of X-axis LRAs and the small control unit with default level-7 intensity (50% duty cycle) and 170 Hz frequency, unless otherwise stated.
@@ 4.1. Latency
Having fine-grained control (DR2) over the actuators in a vibrotactile system means that the data transmission of the system must have low latency and high bandwidth. To determine if the latency of VibraForge was below the human perceivable latency threshold of 45 ms (Vogels, 2004), we measured the end-to-end latency of a PC to a control unit (i.e., the transmission latency of the Bluetooth Low Energy protocol (BLE)), and of a control unit to a vibration unit (i.e., the latency of the custom UART protocol).
To measure the PC to control unit latency, a BLE write GATT command was sent and immediately followed by a BLE read GATT command. A Python timer recorded the write command sending time and read command receiving time and computed the difference as the latency. We repeated this measurement 100 times and found that the average round-trip transmission latency was $26.67\pm 6.53$ ms. Therefore, the one-way transmission latency was approximately $14$ ms. For the control unit to vibration unit latency, repeated measurements using an oscilloscope showed that each command took $125$ us to be transmitted from one unit to the next unit on a chain. Therefore, latency on the longest chain with 16 actuators would accumulate to $2$ ms. Adding the two measurements together would thus result in an overall latency of approximately $16$ ms, which would be virtually imperceptible by humans (Vogels, 2004).
@@ 4.2. Bandwidth
The bandwidth of VibraForge was the number of commands that could be sent and processed per second by control units and vibration units. Bandwidth differs from latency in that asynchronous BLE commands could be continually sent without needing to wait for the previous command to be received (Figure 10). Therefore, bandwidth represented the refresh rate of the vibration commands for each actuator. Since BLE commands (millisecond level) would be processed much slower than UART commands (microsecond level), the bandwidth bottleneck of the system depended on the processing speed of BLE commands on the control unit.
@@
This figure shows the breakdown of data transmission bandwidth and latency, highlighting BLE (Bluetooth Low Energy) transmission and UART communication. The BLE transmission segments are separated by 14 milliseconds, followed by 2 milliseconds of UART communication. The bandwidth is marked as 5 milliseconds or 200 Hz, representing the frequency of data transmission cycles. The diagram visually represents how BLE and UART transmissions alternate in time.
To measure VibraForge’s bandwidth, we wrote a Python testing program to repetitively send standard BLE commands without latency, and another program on the ESP32 (i.e., control unit) to receive the commands and print the receiving time to the serial port. The command processing time was computed as the difference between the receiving time for each command. We sent $1,000$ BLE commands with five UART commands in each of them, and computed an average processing time as $2.96\pm 1.20$ ms. Therefore, for the actual operation of the toolkit, we set the BLE command sending delay to be $5$ ms with some tolerance. This means that the control unit could simultaneously update five vibration units at $200$ Hz refresh rate, satisfying the signal segmentation algorithm requirements.
@@ 4.3. Power Consumption and Battery Life
Power consumption measurements would provide an estimation of how long wearable devices built using VibraForge would function before recharging was necessary (DR6). Power consumption was measured using a bench power supply with a real-time measurement of voltage, current, and power. When only the control unit was connected, it consumed $106$ mA current. When multiple vibration units were connected to the control unit, each vibration unit consumed $2.5$ mA when idle. When a vibration unit was activated, the actuator consumed $150$ mA. Therefore, for a vibrotactile arm display with one control unit and two full chains of vibration units ($16x2)$, the idle current would be $186$ mA and a $500$ mAh LiPo battery would last $2.69$ hours. Even in extreme situations such as multimedia entertainment, running two actuators consistently would consume $486$ mA, leading to $1.03$ hours of operating time. For a larger system with four chains of vibration units ($416x$ 4), the idle current would be $266$ mA, so a $8200$ mAh LiPo battery would last $30.83$ hours. In an extreme situation where eight actuators were continuously activated, the overall current would be $1460$ mA and the battery would last $5.62$ hours.

### v2

@@ 3.2.1. Vibrotactile Actuators
The vibration unit used a square-shaped X-axis LRA (9.5 mm x 9.5 mm, Model L959535, Jinlong Machinery and Electronics (Machinery and Co., 2022)), resonating at 170 Hz with a maximum current of 145 mArms and max acceleration of 0.65±0.12 Grms at 0.9 Vrms. Two additional actuators included a cylindrical Z-axis LRA (8 mm diameter, 3.2 mm height, Model G0832, Jinlong (Machinery and Co., 2022)) resonating at 235 Hz with 0.50±0.08 Grms of acceleration and a maximum current of 90 mArms at 1.8 Vrms, and a VCA (32 mm x 22 mm, Model HD-VA3222, PUI Audio Inc. (PUI Audio, 2023)) with a 80 - 500 Hz response range and 2.52 Gp-p acceleration at 133 Hz, 1.5 Vrms. These actuators covered various vibration directions and intensities to meet the tactile perception needs of different body parts (i.e., spatial acuity (Elsayed et al., 2020) and perceived intensity (Verrillo et al., 1969)). Since they share the same PCB schematics and similar enclosures, we will focus on the X-axis LRA vibration unit in the following sections, with the other units shown in Figure 3.
@@ 3.2.2. Printed Circuit Board Design
Each vibration unit had a PCB driver board to receive UART commands from a control unit and provide fine-grained control over vibration characteristics (DR2). The PCB had a PIC16F18313 MCU, a DRV8837 H-bridge driver, two 4-pin JST-SH headers, and two pogo pins for actuator connections (Figure 4). Upon receiving a UART command from the RX line of the JST connector, the MCU extracted vibration characteristics from the command, generated waveform using an internal PWM generator, and sent it to the actuator through the H-bridge driver. The communication protocol is detailed in Section 3.4.
Figure 4. The circuit schematics and PCB board designs of the vibration units.
@@
The 3D printed enclosures were designed to enclose actuators and PCBs and allow easy attachment to garments (DR4). Compact and lightweight like a coin, they ensured a comfortable wearing experience (DR5). The enclosure included a cap with a screw-nut system to stabilize the PCB, a mounting plate for PCB protection and alignment of the pogo pins, a protective housing with two JST connector holes and heat dissipation vents, and a bottom ring that maintained direct skin contact without sewing (DR4) (Figure 3). The same design strategy was also applied to design enclosures for other actuators.
@@ 3.3. Control Unit
The portable control unit (DR5) featured a PCB extension board with ESP32 MCU, two LiPo batteries, and a 3D-printed enclosure (Figure 5). Multiple versions were designed to accommodate different project requirements, varying in the number of supported chains, the ESP32 model, and battery capacity.
@@
We designed a large control unit supporting up to 8 chains, with 16 units per chain (128 total). Powered by two 3.7 V 8200 mAh LiPo batteries, the unit used a linear regulator for the MCUs and a buck regulator for the actuators, ensuring 5 V stable voltage supplies for both. The extension board had eight JST connectors for vibration units, each linked to a GPIO pin on an Adafruit ESP32 Feather V2 MCU  [^2] https://learn.adafruit.com/adafruit-esp32-feather-v2. The MCU converted Bluetooth commands to UART commands and sent to corresponding vibration units. For devices requiring fewer units, we also created a more compact control unit with two 3.7 V 500 mAh LiPo batteries and an Adafruit QT Py ESP32-S3 MCU  [^3] https://learn.adafruit.com/adafruit-qt-py-esp32-s3, supporting four chain connections.
@@ 3.4. Chain-Connection Method and Data Communication
VibraForge differs from existing toolkits by using a chain-connection method (Figure 6), allowing a single control unit to manage over 100 vibration units (DR1) and simultaneously control multiple actuators’ vibration characteristics (DR2).
@@
Vibration units are connected in chains to receive commands from the ESP32 GPIO pins. We considered two data transmission options (a) data bus protocols (e.g., CAN or I2C) would guarantee minimum latency via broadcasting commands, but require programming different addresses for each unit; (b) serial protocols (e.g., UART) did not require unique address, but passing commands along the chain would increase latency. We decided to use UART because it enabled plug-and-play control without extra programming (DR4) and latency could be reduced to imperceptible levels via high-speed transmission.
Vibration commands were transmitted via UART at 115.2 kHz with parity checks (Figure 7). Each command included two bytes: the first byte encoded the address and start/stop status, and the second byte specified vibration characteristics. Upon receiving the first byte, a unit checked the address, decremented it, and forwarded the command to the next unit. This repeated until the address was zero, thus signaling the target unit had been reached. The targe unit acted based on the start/stop bit: ’0’ (STOP) disabled the actuator, while ’1’ (START) activated it. The second byte detailed the vibration intensity (16 levels from 0% to 100% ) and frequency (i.e., 8 levels following Weber’s factor for frequency differences (Pongrac, 2006), i.e., 123, 145, 170, 200, 235, 275, 322, and 384 Hz).
Figure 7. The two-byte message bits definition and an example of how vibration commands were transmitted along a chain using the ”virtual address” system.
@@ 3.5. Signal Segmentation Algorithm
To render complex waveforms like audio signals (DR2), we implemented a signal segmentation algorithm to convert high-frequency input signals to low-frequency serial commands (Figure 8). For example, a 44100 Hz audio signal with 200 Hz and 5 Hz sine waves would be processed using a short-time Fourier transform (STFT) to extract high-frequency components (main frequency) and a Hilbert transform for low-frequency components (envelope). Both components were then downsampled to 200 Hz and sent to vibration units. Components above 100 Hz would be transmitted via frequencies, and those below 100 Hz as intensities. Technical evaluation (Section 4) confirmed the algorithm ’s reliability in rendering complex waveforms.
@@ 3.6. GUI Editor for Designing Spatialized Patterns
We designed a GUI Editor (Figure 9) to support the creation of spatialized vibration patterns (DR3). The editor combined the waveform editing from single-actuator design software with the canvas design used for multi-actuator design. Its key components were:
@@
  • Waveform Library: The Waveform Library included templates like oscillators and envelopes for creating custom waveforms. Designers could also import JSON waveforms from editors such as Syntacts (Pezent et al., 2020), VIREO (Terenti and Vatavu, 2023), and Meta Haptics Studio (Meta Platforms, 2023a), enabling for the reuse of previous designs (DR3).
  • Waveform Visualizer: The Waveform Visualizer allowed designers to create custom waveforms by combining templates from the Waveform Library via drag-and-drop. It also displayed the composed waveform for selected vibration units during pattern composition.
  • Canvas: The Canvas provides an approximate virtual layout of vibration units and their relative order within each chain. Designers could drag-and-drop a single unit into the Canvas. A ”Create New Chain” function also allowed them to create a new chain of units at once and align the units in grid patterns. Selecting a unit would display its waveform in the Waveform Visualizer.
  • Timeline: The Timeline summarized vibration patterns, with blue blocks indicating activation periods. Moving the slider highlighted active units on the Canvas. Clicking the ”Play” button sent real-time vibration commands to the control unit via Bluetooth.
In a typical spatial pattern design workflow, designers would specify the number of chains and actuators for their system. The GUI then generated a default Canvas with the designated number of actuators, which could be freely adjusted to match their physical layout. Designers then assigned waveforms to each vibration unit, specifying start and end time. Waveforms can be created using templates from the Waveform Library or imported from other software. To review patterns, designers could click on a vibration unit to view its waveform in the Waveform Visualizer or use the Timeline slider to see a high-level summary of vibration patterns over time. This workflow was found to be easy to follow by both experts (Section 5) and novice users (Section 6).
To support the integration of VibraForge with external programming environments (DR3), we created a Python server and a Unity plugin that provided APIs to directly address the VibraForge hardware. The APIs offered functions such as SendCommand(address, start_stop, intensity, frequency) that enabled designers to configure the vibration parameters of a unit without delving into the waveform details, thereby assisting users without extensive technical expertise. The APIs were later used in Case Study 2 and 3 to build VR and robotics applications (Section 5).
@@ 4. Technical Evaluation
In this section, we demonstrate how the VibraForge toolkit fulfills the aforementioned design requirements via various performance metrics. For the evaluations, we used vibration units of X-axis LRAs and the small control unit with a 50% intensity and frequency of 170 Hz, unless otherwise noted.
@@ 4.1. Latency
Low latency and high bandwidth are crucial for data transmission in spatialized vibrotactile systems, particularly in scenarios where user actions trigger immediate tactile feedback, enabling users to respond and adjust their behaviors accordingly. To determine if VibraForge’s latency was below human perceivable threshold of 45 ms (Vogels, 2004), we measured the end-to-end latency of a PC to a control unit (i.e., latency of the Bluetooth Low Energy protocol (BLE)), and of a control unit to a vibration unit (i.e., latency of the UART).
To measure the PC to control unit latency, a BLE GATT write command was sent and immediately followed by a read command. A Python timer computed the difference between the write command sending time and read command receiving time. We repeated this measurement 100 times and found the average round-trip transmission latency was $26.67\pm 6.53$ ms. Therefore, the one-way transmission latency was approximately $14$ ms. For the control unit to vibration unit latency, repeated measurements using an oscilloscope showed that each command took $125$ us to be transmitted from one unit to the next unit. Therefore, latency on a chain with 16 actuators would be $2$ ms. These added to an overall latency of approximately $16$ ms, which would be virtually imperceptible by humans (Vogels, 2004).
@@ 4.2. Bandwidth
VibraForge’s bandwidth was defined as the number of commands that could be sent and processed per second by the control and vibration units. Bandwidth differs from latency in that asynchronous BLE commands could be continually sent without needing to wait for the previous command to be received (Figure 10). Since BLE commands (i.e., millisecond level) would be processed much slower than UART commands (i.e., microsecond level), the bandwidth bottleneck depended on the processing speed of the BLE commands on the control unit.
@@
The breakdown of data transmission bandwidth and latency, highlighting BLE (Bluetooth Low Energy) transmission and UART communication. The BLE transmission segments were separated by 14 milliseconds, followed by 2 milliseconds of UART communication. The bandwidth was 5 milliseconds or 200 Hz, representing the frequency of the data transmission cycles. The diagram represents how BLE and UART transmissions alternate in time.
To measure VibraForge’s bandwidth, a Python testing program repetitively sent BLE commands without latency and another program on the control unit received the commands and printed the receiving time to the serial port. The command processing time was the difference between the receiving time for each command. We sent $1,000$ BLE commands with five UART commands in each of them, and computed an average processing time of $2.96\pm 1.20$ ms. Thus, during actual operation of the toolkit, the BLE command sending delay was set to $5$ ms with some tolerance. The control unit could simultaneously update five vibration units at $200$ Hz refresh rate, satisfying the signal segmentation algorithm requirements.
@@ 4.3. Power Consumption and Battery Life
Power consumption measurements provide an estimation of how long wearable devices would function before recharging was needed (DR5). Power consumption was measured using a bench power supply with a real-time measurement of voltage, current, and power. When only the control unit was connected, it consumed $106$ mA current. When multiple vibration units were connected to the control unit, each vibration unit consumed $2.5$ mA when idle. When a vibration unit was activated, the actuator consumed $150$ mA. Therefore, for a vibrotactile arm display with one control unit and two full chains of vibration units ($16x2)$, the idle current would be $186$ mA and a $500$ mAh LiPo battery would last $2.69$ hours. Even in extreme use cases such as multimedia entertainment, continually running two actuators would consume $486$ mA, leading to $1.03$ hours of operating time. For a larger system with four chains of vibration units ($416x$ 4), the idle current would be $266$ mA, with a $8200$ mAh LiPo battery lasting $30.83$ hours. In an extreme situation with eight actuators continuously activated, the overall current would be $1460$ mA and the battery would last $5.62$ hours.